# Finite–slope universal Rankin–Selberg $p$-adic $L$-function

### v1

@@ 2. Target main theorem (precise special case)
In his work on $p$-adic $L$-functions in universal deformation families, Loeffler formulates a general conjecture (Conjecture 2.8 of [13]) predicting the existence and interpolation properties of rank 0 Euler systems (i.e. $p$-adic $L$-functions) attached to 0-Panchishkin families $(R,V,V^{+})$ over suitable “big parabolic eigenvarieties”. His main examples include the half–ordinary Rankin–Selberg tensor considered in Example 3.17 of loc. cit., which is precisely the global deformation setting adopted in Step 1 of this article.
@@
In the ordinary case, this conjectural picture has now been confirmed for the Rankin–Selberg tensor by the recent work of Hao–Loeffler [15]. Building on Urban’s nearly overconvergent three–variable Rankin–Selberg $p$-adic $L$-function [7] and the Beilinson–Flach Euler system of Loeffler–Zerbes [21], they construct a universal Rankin–Selberg $p$-adic $L$-function for an ordinary Hida family tensored with a universal deformation family, and prove that it interpolates all Deligne–critical Rankin–Selberg values in this setting, together with a functional equation.
@@
However, both [13] and [15] work under a global ordinarity (or nearly–ordinarity) hypothesis at $p$, and the $p$-adic $L$-functions they construct live naturally over ordinary eigenvarieties or ordinary loci in universal deformation spaces. In contrast, in the present paper we fix a classical point $x_{0}=(f_{0},g_{0})$ at which the $f_{0}$–factor is of non–ordinary finite slope and we work in a neighbourhood $U$ of $x_{0}$ in a parabolic eigenvariety in the sense of Barrera Salazar–Williams [9]. Theorem 3 should therefore be viewed as a first verification, in a concrete $\operatorname{GL}_{2}\times\operatorname{GL}_{2}$ situation, of the finite–slope variants of Loeffler’s Conjecture 2.8 over big parabolic eigenvarieties suggested in [13, §5.5, §6].
3.2. Comparison with existing finite–slope constructions.
@@
Finite–slope $p$-adic $L$-functions attached to families of automorphic forms have been constructed in several other settings, but always in a way that is complementary to the present work.
@@
  • For Rankin–Selberg convolutions of modular forms, Urban’s three–variable $p$-adic $L$-function is constructed on eigenvarieties using nearly overconvergent modular forms and overconvergent cohomology [7], and requires a nearly ordinary hypothesis at $p$. These functions live over eigenvarieties and are not formulated in terms of universal Galois deformation spaces.
@@
  • The geometric methods of Andreatta–Iovita and their collaborators yield triple product $p$-adic $L$-functions for finite slope families of modular forms via the theory of overconvergent sheaves and the spectral halo [11]. Again, the base spaces are eigenvarieties, and the Galois representations are not packaged as a single Panchishkin family $(R,V,V^{+})$ over a universal deformation ring.
@@
  • Barrera Salazar–Dimitrov–Williams construct finite–slope $p$-adic $L$-functions for Shalika families on $\operatorname{GL}_{2n}$ over parabolic eigenvarieties, and relate their existence to the local geometry of these eigenvarieties [12]. Their methods are adapted to the standard representation of $\operatorname{GL}_{2n}$ and do not provide a universal deformation interpretation for Rankin–Selberg tensors of $\operatorname{GL}_{2}\times\operatorname{GL}_{2}$.
By contrast, the present article keeps the universal deformation viewpoint of [13, 15], starting from the half–ordinary Panchishkin family $(R,V,V^{+})$ of Example 3.17 of [13], and then passes to a finite–slope neighbourhood $U$ on the parabolic eigenvariety through the identification $U\simeq U_{R}\subset\operatorname{Spf}(R)_{\mathrm{rig}}$ of Step 1.4. The resulting finite–slope universal Rankin–Selberg $p$-adic $L$-function
$L_{p}^{\mathrm{fs}}\;=\;\mathcal{L}_{V_{U},V_{U}^{+}}(\mathcal{BF}_{U})\in\mathcal{O}(U\times\mathcal{W})$
@@
simultaneously interpolates the critical Rankin–Selberg values for all classical points in $U$ and specialises to the ordinary universal Rankin–Selberg $p$-adic $L$-function at ordinary points. In this way it provides a conceptual bridge between the ordinary universal results of [13, 15] and the finite–slope eigenvariety constructions along the lines of Urban and Andreatta–Iovita [7, 11].
3.3. Why the finite–slope case is difficult.
@@
From the perspective of [13], it is natural to expect that a finite–slope analogue of Conjecture 2.8 should hold whenever one can attach a $(\varphi,\Gamma)$-module to the Panchishkin family $(R,V,V^{+})$ and construct an appropriate Perrin–Riou regulator [18]. However, there are several serious obstacles which have so far prevented a general theory, and which Theorem 3 overcomes in the specific Rankin–Selberg case treated here:
@@
  • Cohomology and base change. Even after triangulation, one must control the $(\varphi,\Gamma)$- and Iwasawa cohomology of $D$ and $D^{+}$ in families, with good base–change and perfectness properties, in order to set up a Perrin–Riou type regulator over the affinoid algebra $\mathcal{O}(U)$. This is achieved here by systematically using the cohomological machinery of Kedlaya–Pottharst–Xiao [19], which guarantees that the relevant cohomology complexes lie in the perfect derived category and behave well under specialisation.
@@
  • Family Perrin–Riou regulator for a genuine finite–slope Panchishkin family. Existing constructions of big logarithms and regulators in families (following Perrin–Riou, Coleman, Loeffler–Zerbes, Nakamura, Pottharst, etc.; see for instance [18, 21, 16]) either treat a single de Rham representation or work under an ordinarity hypothesis. The map
    $\mathcal{L}_{V_{U},V_{U}^{+}}:H^{1}_{\mathrm{Iw}}(\mathbb{Q}_{p},V_{U}^{\ast}(1))\;\longrightarrow\;\mathcal{H}(\Gamma)\,\widehat{\otimes}_{\mathbb{Q}_{p}}\mathcal{O}(U)$
@@
    constructed in Step 3 is, to the best of our knowledge, the first instance of a Perrin–Riou regulator attached to a genuinely finite–slope Panchishkin family over a higher–dimensional parabolic eigenvariety in the universal deformation setting.
@@
  • Compatibility with Beilinson–Flach Euler systems. Finally, one must show that the universal Beilinson–Flach classes of Loeffler–Zerbes [21], interpreted in the universal deformation setting (following [15]), specialise correctly along $U$ and lie in the critical Iwasawa cohomology group on which $\mathcal{L}_{V_{U},V_{U}^{+}}$ is defined. This compatibility is crucial in order to define $L_{p}^{\mathrm{fs}}$ and deduce its interpolation formula from the explicit reciprocity laws for Beilinson–Flach elements.
These issues explain why, despite the conceptual clarity of the finite–slope variant of Conjecture 2.8, no general finite–slope universal Rankin–Selberg $p$-adic $L$-function had been constructed before.
3.4. Future directions.
@@
The techniques developed in this article suggest several natural directions for further work:
@@
  • Beyond the small–slope, non–critical range. A first goal would be to relax the crystalline small–slope and non–criticality hypotheses at $p$ imposed in Theorem 3. This would require a finer analysis of triangulations and local Galois representations at critical slope, possibly using the spectral–halo techniques of Andreatta–Iovita and the overconvergent cohomology methods of Barrera Salazar–Williams [11, 9].
@@
  • Other groups and higher rank. It should be possible, at least in principle, to adapt the strategy of Steps 1–3 to other instances of Loeffler’s conjectural framework, for example to Shalika families on $\operatorname{GL}_{2n}$ or to $\mathrm{GSp}_{4}$-type Rankin–Selberg convolutions, combining the parabolic eigenvarieties of Barrera Salazar–Williams [9, 12] with the Euler systems and Perrin–Riou regulators available in those settings.
  • Iwasawa–theoretic applications and main conjectures. The finite–slope universal Rankin–Selberg $p$-adic $L$-function $L_{p}^{\mathrm{fs}}$ should control Selmer groups for the specialisations $V_{x}$ in a manner predicted by the Bloch–Kato conjectures and noncommutative Iwasawa theory. One long–term objective is to formulate and prove main conjectures relating characteristic ideals of universal Selmer modules over $R$ to the ideal generated by $L_{p}^{\mathrm{fs}}$.
  • Functional equations and exceptional zeros. The universal functional equation obtained in [15] for the ordinary Rankin–Selberg $p$-adic $L$-function suggests that an analogue should exist for $L_{p}^{\mathrm{fs}}$, relating its values (or derivatives) at critical points to those of other $p$-adic $L$-functions. Understanding such a functional equation in the finite–slope setting could lead to new results on exceptional zeros and on the variation of local epsilon factors in families.
  • Refined arithmetic information. Finally, one may hope to extract finer arithmetic invariants from $L_{p}^{\mathrm{fs}}$, such as $p$-adic heights or regulators on motivic cohomology classes varying in families, in the spirit of the work of Darmon–Rotger and Hsieh on triple product $p$-adic $L$-functions. This would provide further evidence that big parabolic eigenvarieties are indeed the correct parameter spaces for Euler systems and $p$-adic $L$-functions beyond the ordinary setting.
In summary, Theorem 3 provides the first finite–slope verification of Loeffler’s universal conjectural picture in the Rankin–Selberg case, and the methods developed here should be of independent interest for the study of $p$-adic $L$-functions and Euler systems on parabolic eigenvarieties beyond the ordinary locus.
@@ 2 Proof and Results
Step 1. Global deformation space and half–ordinary Panchishkin family
@@
arising from normalised cuspidal Hecke eigenforms $f_{0}$ and $g_{0}$ of tame level $1$ (or more generally of fixed tame level prime to $p$), with $\bar{\rho}_{1}$ ordinary at $p$ and $\bar{\rho}_{2}$ subject to the usual Taylor–Wiles hypotheses. We write $\mathcal{O}$ for a finite extension of $\mathbb{Z}_{p}$ containing the Hecke eigenvalues of $f_{0}$ and $g_{0}$, and we let $\mathrm{CNL}_{\mathcal{O}}$ denote the category of complete Noetherian local $\mathcal{O}$–algebras with residue field $\mathbb{F}$.
(1.1) Ordinary universal deformation ring for $\bar{\rho}_{1}$.
Let $R^{\mathrm{ord}}(\bar{\rho}_{1})$ denote the universal deformation ring parametrising ordinary deformations of $\bar{\rho}_{1}$ as a $G_{\mathbb{Q},\{p\}}$–representation in the sense of Hida and Mazur: for any $A\in\mathrm{CNL}_{\mathcal{O}}$, a deformation
@@
$0\longrightarrow V_{1,A}^{\mathrm{ord},+}\longrightarrow V_{1,A}^{\mathrm{ord}}\big|_{G_{\mathbb{Q}_{p}}}\longrightarrow V_{1,A}^{\mathrm{ord},-}\longrightarrow 0$
where $V_{1,A}^{\mathrm{ord},+}$ is a rank–$1$ direct summand on which $G_{\mathbb{Q}_{p}}$ acts by an unramified character lifting the pp–adic unit root of the Hecke polynomial of $f_{0}$.
By Mazur’s deformation theory together with Hida theory and the work of Böckle and Emerton (see, for example, [13, §3.2, Prop. 3.14 and Thm. 3.4]), this ordinary deformation problem is representable by a complete Noetherian local $\mathcal{O}$–algebra $R^{\mathrm{ord}}(\bar{\rho}_{1})$, and there exists a universal ordinary deformation
$\rho_{1}^{\mathrm{ord}}:G_{\mathbb{Q},\{p\}}\longrightarrow\operatorname{GL}_{2}(R^{\mathrm{ord}}(\bar{\rho}_{1}))$
with underlying free rank–$2$ $R^{\mathrm{ord}}(\bar{\rho}_{1})$–module $V_{1}^{\mathrm{ord}}$.
We denote by $V_{1}^{\mathrm{ord},+}\subset V_{1}^{\mathrm{ord}}\big|_{G_{\mathbb{Q}_{p}}}$ the rank–$1$ direct summand giving the ordinary filtration at $p$, and by $V_{1}^{\mathrm{ord},-}$ the corresponding quotient. The ring $R^{\mathrm{ord}}(\bar{\rho}_{1})$ is finite flat over the weight space and is canonically isomorphic to the localised ordinary Hecke algebra $T^{\mathrm{ord}}(\bar{\rho}_{1})$ acting on ordinary pp–adic modular forms of tame level $1$; in particular $R^{\mathrm{ord}}(\bar{\rho}_{1})$ is reduced and equidimensional of relative dimension $1$ over $\mathcal{O}$ (see [13, §3.2, Prop. 3.14]).
For brevity we set
$R_{1}:=R^{\mathrm{ord}}(\bar{\rho}_{1}),\qquad V_{1}^{\mathrm{ord}}\text{ and }V_{1}^{\mathrm{ord},+}\text{ as above.}$
(1.2) The unrestricted universal deformation ring and representation for $\bar{\rho}_{2}$.
For the second factor, we do not impose any ordinary or nearly–ordinary local condition at $p$. We let $R(\bar{\rho}_{2})$ be the universal deformation ring parametrising deformations of $\bar{\rho}_{2}$ as a $G_{\mathbb{Q},\{p\}}$–representation (i.e. unramified outside $p$), as in [13, Def. 3.3]. Thus, for any $A\in\mathrm{CNL}_{\mathcal{O}}$, a deformation to $A$ is simply a continuous representation

### v2

@@ 1.3 Discussion and relation to previous work
In his work on $p$-adic $L$-functions in universal deformation families, Loeffler formulates in [13, Conj. 2.8] a general conjecture predicting the existence and interpolation properties of rank 0 Euler systems (equivalently, $p$-adic $L$-functions) attached to 0-Panchishkin families $(R,V,V^{+})$ over suitable “big parabolic eigenvarieties”. A main example is the half–ordinary Rankin–Selberg tensor considered in [13, Ex. 3.17], which is precisely the global deformation setting considered here.
@@
In the ordinary case, this conjectural picture has now been confirmed for the Rankin–Selberg tensor by recent work of Hao–Loeffler [15]. Building on Urban’s nearly overconvergent three–variable Rankin–Selberg $p$-adic $L$-function [7] and the Beilinson–Flach Euler system of Loeffler–Zerbes [21], they construct a universal Rankin–Selberg $p$-adic $L$-function for an ordinary Hida family tensored with a universal deformation family and prove that it interpolates all Deligne–critical Rankin–Selberg values in this setting, together with a functional equation. Their construction is analytic and does not rely on the existence of a universal Euler system.
@@
By contrast, in this work we fix a classical point $x_{0}=(f_{0},g_{0})$ at which the $f_{0}$-factor has non–ordinary finite slope and we work in a neighbourhood $U$ of $x_{0}$ in a parabolic eigenvariety in the sense of Barrera Salazar–Williams [9]. Theorem 1.4 should therefore be viewed as a conditional finite-slope analogue of the Rankin–Selberg case of Loeffler’s conjecture over big parabolic eigenvarieties, in a concrete $\operatorname{GL}_{2}\times\operatorname{GL}_{2}$ setting.
Comparison with existing finite-slope constructions
@@
Finite-slope $p$-adic $L$-functions attached to families of automorphic forms have been constructed in several other settings, but usually not in the universal deformation framework.
@@
For Rankin–Selberg convolutions of modular forms, Urban’s three–variable $p$-adic $L$-function [7] is constructed on eigenvarieties using nearly overconvergent modular forms and overconvergent cohomology and requires a nearly ordinary hypothesis at $p$. These functions live over eigenvarieties and are not formulated in terms of universal Galois deformation spaces.
@@
The geometric methods of Andreatta–Iovita and their collaborators yield triple product $p$-adic $L$-functions for finite-slope families of modular forms via the theory of overconvergent sheaves and the spectral halo [11]. Here again the base spaces are eigenvarieties and the Galois representations are not packaged as a single Panchishkin family over a universal deformation ring.
@@
Barrera Salazar–Dimitrov–Williams construct finite–slope $p$-adic $L$-functions for Shalika families on $\operatorname{GL}_{2n}$ over parabolic eigenvarieties and relate their existence to the local geometry of these eigenvarieties [12]. Their methods are adapted to the standard representation of $\operatorname{GL}_{2n}$ and do not provide a universal deformation interpretation for Rankin–Selberg tensors of $\operatorname{GL}_{2}\times\operatorname{GL}_{2}$.
@@
In contrast, the present article maintains the universal deformation viewpoint of [13, 15], starting from the half–ordinary Panchishkin family $(R,V,V^{+})$ of [13, Ex. 3.17], and then passing to a finite-slope neighbourhood $U$ on the parabolic eigenvariety through the identification of an ordinary locus in $U$ with an affinoid subdomain of $\operatorname{Spf}(R)^{\operatorname{rig}}$. The function $L_{p}^{\mathrm{fs}}$ constructed under Conjecture 2.15 simultaneously interpolates the critical Rankin–Selberg values for all classical points in $U$ and specialises, at ordinary points, to the ordinary universal Rankin–Selberg $p$-adic $L$-function of [15]. In this way it provides a conceptual bridge between the ordinary universal results and finite-slope eigenvariety constructions in the spirit of Urban and Andreatta–Iovita.
Why the finite-slope case is difficult
@@
From the perspective of [13], it is natural to expect that a finite-slope analogue of Conjecture 2.8 should hold whenever one can attach a $(\varphi,\Gamma)$-module to the local Galois representation in a family and construct an appropriate Perrin–Riou regulator. However, there are several substantial obstacles which have so far prevented a general theory, and which in the Rankin–Selberg case are addressed here.
@@
Second, even after triangulation, one must control the $(\varphi,\Gamma)$- and Iwasawa cohomology of $D_{U}$ and $D_{U}^{+}$ in families, with good base-change and perfectness properties, in order to set up a Perrin–Riou type regulator over the affinoid algebra $\mathcal{O}(U)$. This is achieved here by systematically using the cohomological machinery of Kedlaya–Pottharst–Xiao [19], which guarantees that the relevant cohomology complexes lie in the perfect derived category and behave well under specialisation.
@@
Third, existing constructions of big logarithms and regulators in families (following Perrin–Riou, Coleman, Loeffler–Zerbes, Nakamura, Pottharst, and others; see for instance [18, 21, 16]) typically treat a single de Rham representation or work under an ordinarity hypothesis. The map
$\mathcal{L}_{V_{U},V_{U}^{+}}:H^{1}_{\operatorname{Iw}}(\mathbb{Q}_{p},V_{U}^{\ast}(1))\longrightarrow\mathcal{H}(\Gamma)\widehat{\otimes}_{\mathbb{Q}_{p}}\mathcal{O}(U)$
@@
constructed in this paper is, to the author’s knowledge, one of the first instances of a Perrin–Riou regulator attached to a genuinely finite-slope Panchishkin family over a higher-dimensional parabolic eigenvariety in the universal deformation setting.
@@
Finally, to obtain a $p$-adic $L$-function one needs an Euler system which specialises correctly along $U$ and lies in the critical Iwasawa cohomology group on which $\mathcal{L}_{V_{U},V_{U}^{+}}$ is defined. This compatibility is crucial in order to define $L_{p}^{\mathrm{fs}}$ as $\mathcal{L}_{V_{U},V_{U}^{+}}(\mathcal{BF}_{U})$ and deduce its interpolation formula from explicit reciprocity laws for Beilinson–Flach elements. At present the existence of such a global class $\mathcal{BF}_{U}$ for the half–ordinary universal deformation family is conjectural (Conjecture 2.15); in the ordinary setting [15] construct the universal Rankin–Selberg $p$-adic $L$-function analytically, without passing through a universal Euler system.
Future directions
@@
The techniques developed here suggest several directions for further work, which we only mention briefly.
@@
One would like to relax the crystalline small-slope and non-criticality hypotheses at $p$. This would require a finer analysis of triangulations and local Galois representations at critical slope, possibly using spectral-halo techniques and overconvergent cohomology as in [11, 9].
@@
It should be possible to adapt the strategy of the later sections to other instances of Loeffler’s conjectural framework, for example to Shalika families on $\operatorname{GL}_{2n}$ or to $\operatorname{GSp}_{4}$-type Rankin–Selberg convolutions, combining parabolic eigenvarieties [9, 12] with existing Euler systems and regulators in those settings.
Assuming the Euler-system conjecture, it would be natural to investigate Iwasawa-theoretic applications and formulate main conjectures relating characteristic ideals of universal Selmer modules over $R$ to the ideal generated by $L_{p}^{\mathrm{fs}}$.
Finally, one expects a functional equation for $L_{p}^{\mathrm{fs}}$ relating its values (or derivatives) at critical points to those of other $p$-adic $L$-functions, and one might hope to extract finer arithmetic invariants such as $p$-adic heights and regulators on motivic cohomology classes varying in families.
@@ 2 Proof and results
In this section we explain how the hypotheses (H1)–(H3) lead to the construction of the family regulator and, under Conjecture 2.15, to the $p$-adic $L$-function $L_{p}^{\mathrm{fs}}$ of Theorem 1.4. The discussion is divided into three steps.
2.1 Step 1: Global deformation space and half–ordinary Panchishkin family
@@
arising from normalised cuspidal Hecke eigenforms $f_{0}$ and $g_{0}$ of tame level $1$ (or more generally of fixed tame level prime to $p$), with $\bar{\rho}_{1}$ ordinary at $p$ and $\bar{\rho}_{2}$ subject to the usual Taylor–Wiles hypotheses. We write $\mathcal{O}$ for a finite extension of $\mathbb{Z}_{p}$ containing the Hecke eigenvalues of $f_{0}$ and $g_{0}$, and we let $\mathrm{CNL}_{\mathcal{O}}$ denote the category of complete Noetherian local $\mathcal{O}$-algebras with residue field $\mathbb{F}$.
2.1.1 Ordinary universal deformation ring for $\bar{\rho}_{1}$
Let $R^{\mathrm{ord}}(\bar{\rho}_{1})$ denote the universal deformation ring parametrising ordinary deformations of $\bar{\rho}_{1}$ as a $G_{\mathbb{Q},\{p\}}$-representation in the sense of Hida and Mazur: for any $A\in\mathrm{CNL}_{\mathcal{O}}$, a deformation
@@ 2.1.1 Ordinary universal deformation ring for $\bar{\rho}_{1}$
$0\longrightarrow V^{\mathrm{ord},+}_{1,A}\longrightarrow V^{\mathrm{ord}}_{1,A}\big|_{G_{\mathbb{Q}_{p}}}\longrightarrow V^{\mathrm{ord},-}_{1,A}\longrightarrow 0$
where $V^{\mathrm{ord},+}_{1,A}$ is a rank one direct summand on which $G_{\mathbb{Q}_{p}}$ acts by an unramified character lifting the $p$-adic unit root of the Hecke polynomial of $f_{0}$.
By Mazur’s deformation theory together with Hida theory and the work of Böckle and Emerton (see for example [13, §3.2, Prop. 3.14 and Thm. 3.4]), this ordinary deformation problem is representable by a complete Noetherian local $\mathcal{O}$-algebra $R^{\mathrm{ord}}(\bar{\rho}_{1})$, and there exists a universal ordinary deformation
$\rho_{1}^{\mathrm{ord}}:G_{\mathbb{Q},\{p\}}\longrightarrow\operatorname{GL}_{2}\big(R^{\mathrm{ord}}(\bar{\rho}_{1})\big)$
with underlying free rank two $R^{\mathrm{ord}}(\bar{\rho}_{1})$-module $V_{1}^{\mathrm{ord}}$. We denote by $V_{1}^{\mathrm{ord},+}\subset V_{1}^{\mathrm{ord}}\big|_{G_{\mathbb{Q}_{p}}}$ the rank one direct summand giving the ordinary filtration at $p$, and by $V_{1}^{\mathrm{ord},-}$ the corresponding quotient.
The ring $R^{\mathrm{ord}}(\bar{\rho}_{1})$ is finite flat over the weight space and is canonically isomorphic to the localised ordinary Hecke algebra $T^{\mathrm{ord}}(\bar{\rho}_{1})$ acting on ordinary $p$-adic modular forms of tame level $1$; in particular $R^{\mathrm{ord}}(\bar{\rho}_{1})$ is reduced and equidimensional of relative dimension $1$ over $\mathcal{O}$ (see [13, §3.2, Prop. 3.14]). For brevity we set
$R_{1}:=R^{\mathrm{ord}}(\bar{\rho}_{1}).$
2.1.2 The unrestricted universal deformation ring and representation for $\bar{\rho}_{2}$
For the second factor, we do not impose any ordinary or nearly ordinary local condition at $p$. We let $R(\bar{\rho}_{2})$ be the universal deformation ring parametrising deformations of $\bar{\rho}_{2}$ as a $G_{\mathbb{Q},\{p\}}$-representation (unramified outside $p$), as in [13, Def. 3.3]. Thus, for any $A\in\mathrm{CNL}_{\mathcal{O}}$, a deformation to $A$ is simply a continuous representation